# CLIMB: A Benchmark of Clinical Bias in Large Language Models

### v1

@@ 3 CLIMB
We focus on the bias associated and reflected in LLMs’ diagnostic capability, where the diagnoses are identified by the International Classification of Diseases, Tenth Revision, Clinical Modification (ICD-10-CM) codes icd (2023). These well-known codes are adopted by numerous healthcare providers to catalog all medical diagnoses. It also categorizes diagnoses from high (Level 1) to low (Level 5) levels: L1 - Chapters; L2 - Blocks; L3 - Categories; L4 - Sub-categories; L5 - Full code.
@@
Diagnostic Task and Dataset
Intrinsic bias studies the associations between demographics and diagnosis in LLMs. We adopt people’s names as a proxy for different demographics as existing literature has done Sandoval et al. (2023); An et al. (2023) and obtain the diagnoses from ICD-10-CM. Extrinsic bias is defined as the change in model’s performance when we switch the demographic information in a clinical note where we use the diagnosis task introduced in CliBench Ma et al. (2024) as the end task. CliBench contains representative clinical cases induced from the Medical Information Mart for Intensive Care (MIMIC)-IV database, which is comprised of de-identified Electronic Health Records (EHR) for patients admitted to the Beth Israel Deaconess Medical Center. [^3] We adopt the same pre-processing as in CliBench. Unlike previous medical QA datasets, which have limited diagnosis options, the patient instances we use here are real clinical cases with a large number of diagnosis options. This allows us to test LLMs’ biases in realistic clinical settings. To evaluate the extrinsic bias, based on the expert-verified patient’s medical records, we instruct an LLM to provide multiple diagnosis predictions from a list of 94,739 diseases.
@@
In the medical domain, it is an important and essential issue to recognize that certain diagnoses are specific to particular demographics. This specificity arises from correct causal factors that affect different groups in varying ways. Properly acknowledging this distinction ensures accurate and fair clinical assessments, thereby improving healthcare outcomes for all demographic groups. We utilize the female-only [^4] https://www.icd10data.com/ICD10CM/Codes/Rules/Female_Diagnosis_Codes and male-only [^5] https://www.icd10data.com/ICD10CM/Codes/Rules/Male_Diagnosis_Codes codes defined in ICD-10-CM to identify sex-specific diagnoses, resulting in 3,116 and 529 diagnoses respectively. In our extrinsic bias evaluation dataset, 23.6% of the records contain at least one sex-specific diagnosis. We consider these records as sex-specific and others as sex-neutral. Conversely, identifying ethnicity-specific diagnoses is challenging because the absence of robust clinical evidence linking specific diagnoses to particular ethnic groups, without resorting to stereotypes or problematic assumptions, complicates the establishment of such categories Omiye et al. (2023). Therefore, we leave it as future work and only focus on the sex-specific diagnoses in this work.
@@ 3.1 Evaluation of Intrinsic Bias
In psychological research, Implicit Association Test (IAT) is a popular tool used to reveal the unconscious bias of humans by measuring the strength of subjects’ automatic associations between concepts and targets Greenwald et al. (1998). In the context of NLP, IAT has been adapted to assess the intrinsic bias of language models through representation associations between target words (e.g., names of different genders) and attribute words (e.g., terms related to careers or families) that are embedded in models’ internal representation Caliskan et al. (2016); Guo and Caliskan (2020); Bai et al. (2024). Our evaluation of intrinsic bias is also based on IAT with several important modifications.
To adapt IAT for our diagnostic task, we use diagnoses and demographic groups as attributes and targets respectively. We use descriptions from the ICD-10-CM as stimuli of diagnoses and names as stimuli of both sex and ethnicity demographic information following traditional IAT Greenwald et al. (1998); Caliskan et al. (2016). However, existing IAT only provides 2 ethnicity groups (“White” and “Black”) based on names. Instead we obtain fine-grained ethnicity categorizations by collecting the top five most popular names by sex (Female and Male) and ethnicity (White non-Hispanic, Black non-Hispanic, Hispanic and Asian and Pacific Islander)  [^6] We call them as “White”, ”Black”, ”Hispanic” and ”Asian” for simplicity, which is on a par with extrinsic part. from the “Popular Baby Names” dataset  [^7] https://data.cityofnewyork.us/Health/Popular-Baby-Names/25th-nujf, which is a typical way adopted in existing works such as An et al. (2023) and Sandoval et al. (2023).
@@
$\text{MAD}=\frac{1}{n}\sum_{i=1}^{n}|x_{i}-\mu|\vspace{-0.5em}$ (1)
In embedding-based IATs such as WEAT Caliskan et al. (2016) and CEAT Guo and Caliskan (2020), association scores are calculated using embedding distances (e.g., cosine similarity) between attributes and targets that are not applicable in recent LLMs. Instead, for decoder-only LLMs that are designed for autoregressive generation, we follow an alternative approach using next-token probability Chen et al. (2024b); Dong et al. (2023) for the association calculation. Specifically, we adopt the prompt “{$d_{i}$} is related to the name:” and measure the probability of each name ($n_{i}$) appearing as the next token for the association score corresponding to diagnosis $d_{i}$. [^8] The intuition is that this template clearly instructs an LLM to assign names based on diagnosis, while keeping the model unconscious about the specific bias we are investigating. For names tokenized into multiple sub-tokens, we calculate the joint probability of all sub-tokens. With the aforementioned approach, given a set of diagnoses $D=\{d_{1},d_{2},d_{3},\ldots\}$ and a set of demographic groups $G=\{g_{1},g_{2},g_{3},\ldots\}$, the association score $s\left(d_{i},g_{j}\right)$ is defined as:
$s\left(d_{i},g_{j}\right)=\frac{1}{\left|W_{g_{j}}\right|}\sum_{k=1}^{\left|W_{g_{j}}\right|}p\left(w_{k}^{g_{j}}|d_{i}\right)\vspace{-0.5em}$ (2)
@@
$\text{\mbox{AssocMAD}}_{d_{i}}=\frac{1}{|G|}\sum_{j=1}^{|G|}\left|s\left(d_{i},g_{j}\right)-\mu_{s\left(d_{i}\right)}\right|\vspace{-0.5em}$ (3)
@@
$\text{\mbox{AssocMAD}}_{d_{i}}=\frac{1}{|G|}\sum_{j=1}^{|G|}\left|\frac{s\left(d_{i},g_{j}\right)-\mu_{s\left(d_{i}\right)}}{\mu_{s\left(d_{i}\right)}}\right|\vspace{-0.5em}$ (4)
@@
$\qquad\qquad\end{multlined}\text{Single-demographic AssocMAD}_{d_{i},x}\qquad\qquad\\ =\frac{1}{|G_{\text{x}}|}\sum_{j=1}^{|G_{\text{x}}|}\left|\frac{s_{\text{x}}\left(d_{i},g_{j}^{\text{x}}\right)-\mu_{s_{\text{x}}\left(d_{i}\right)}}{\mu_{s_{\text{x}}\left(d_{i}\right)}}\right|\qquad\qquad\vspace{-0.5em}$ (5)
@@ 3.2 Evaluation of Extrinsic Bias
Extrinsic bias aims to reflect the disparity of models’ performance on end tasks towards various demographics. To set up a controlled setting and reflect the influence of the bias attribute only, we measure the extrinsic bias of LLMs by their output difference when given factual and counterfactual demographic input prompts.
From the evaluation set of Ma et al. (2024), which consists of clinical cases balanced across chapters of diseases, service departments, and care units, we use all 199 instances with “white” ethnicity, “male” sex, and “other” insurance type 999“Medicare” and “Medicaid” are insurance programs that aim at helping minorities. “Other” covers a more privileged group of people, which is more suitable to serve as a baseline in our experiment. , which is the most privileged group within this dataset, as the default diagnosis data instances. In Appendix A, we provide data statistics of this curated evaluation set compared to the original one in CliBench. For more details about the data details, please refer to the paper of CliBench.
For each sampled admission record in our evaluation set, we follow the same prompt template introduced in CliBench including the clinical notes and demographic information. A full example of the prompt template can be found in the paper of CliBench. To build the counterfactual example, we replace the demographic information with a corresponding group. When the demographic information is about sex, we also replace pronouns and gendered honorifics in the instance to ensure the context remains coherent and accurate. We define extrinsic bias score as the difference in LLM’s performance on the diagnosis prediction task. In real-world clinical settings, we prioritize reducing Type 2 errors to avoid missing diagnoses. To achieve this, we use the difference in recall rate averaged on diagnostic abstraction levels while inputting default and counterfactual patient information to the models. To mitigate the potential effect of positional bias of demographic information, we computed the average score of putting demographic information before and after clinical notes. For the diagnosis prediction task, we follow CliBench to extract valid diagnosis predictions from natural language output generated by LLMs and convert the natural language to ICD code whose description has the highest cosine similarity with the text.
@@ 4.2.1 Sex-neutral Diagnoses
Table 1 reports the intrinsic bias for sex-neutral diagnoses in terms of models’ AssocMAD, which measure the disparity across all demographic groups. A smaller value indicates a smaller disparity (i.e., less biased). We summarize several key observations from this scenario: I. Overall, all models show smaller AssocMAD at a higher level, which can be attributed to the very broad coverage of diagnoses in coarser analyses. On average, Mistral Instruct v0.2 7B and Meditron 7B (row #2 and #7) perform the best with the lowest AssocMAD. Mistral Instruct v0.2 7B achieves the best at higher levels (L1 to L3), whereas Meditron 7B achieves the best at lower levels (L4 to L5). LLaMA2 Chat 13B (#9) performs the worst across all levels, indicating the most biased. II. Interestingly, larger models such as Mistral Instruct v0.1 46.7B (#5), LLaMA2 Chat 13B (#9), and LLaMA3 Instruct 70B (#13) did not outperform their smaller 7B versions (#1, 6 and 11), which suggests that simply increasing model size does not necessarily mitigate bias and may exacerbate it due to the model’s complexity and the varied data it is exposed to during training. For the most up-to-date version of Mistral and LLaMA, Mistral v0.3 7B (#3) is more biased than LLaMA3 Instruct 8B (#11) with an averaged AssocMAD of 0.41, compared to 0.36. III. Among medically adapted models, Meditron (7B) stands out, as introduced earlier, with the lowest AssocMAD. However, with the same based model (i.e., LLaMA2 Chat 7B; #6), Asclepius 7B (#8) achieves slighter improvement. A possible reason may be that its fine-tuning data (i.e., PMC-Patients case reports) contain more relations between the diagnosis codes and patients’ demographics. IV. Models pre-trained or fine-tuned on medical corpora with incorrect association may even introduce more intrinsic bias. For example, both BioMistral DARE 7B (#4) and OpenBioLLM 8B (#12) are worse than their base models Mistral Instruct v0.1 7B (#1) and LLaMA3 Instruct 8B (#11), respectively.
@@ 4.2.2 Sex-specific Diagnoses
We report correctness of sex preference for sex-specific diagnoses in Table 3. Although some models demonstrate high correctness for either female-only or male-only diagnoses, they often fail to exhibit a balanced and reasonable sex preference for both. For instance, LLaMA2 Chat 13B shows a high correctness of 0.94 for female-only diagnoses but a low correctness of 0.14 for male-only diagnoses. This indicates it still prefers females even for male-only diagnoses. Similarly, LLaMA3 Instruct 8B shows a striking imbalance with correctness of 0.80 for female-only and 0.03 for male-only diagnoses. This trend is also shown in the medically adapted models and is even more severe than the corresponding base version (e.g., Asclepius compared to LLaMA2 Chat 7B), which suggests that the medically adapted training may not help the model learn the correct knowledge of sex preference for these sex-specific diagnoses.
@@ 4.3.1 Overall Result
We expect a model could be slightly sensitive to sex and ethnicity as they are potential causal factors to diagnosis, but non-sensitive to insurance. From Table 4, we observe:
Decreased diagnosis performance for female instances across models.
Comparing models when given factual and counterfactual sex information, we observed a consistent performance drop in almost all models when we provide counterfactual female sex information. However, when given counterfactual ethnicity or insurance, the results are mixed.
Medically adapted LLMs are more sensitive to demographic information change
Comparing the general domain models and their corresponding medically adapted models, we observed a more significant performance change in medically adapted ones, indicating they are more sensitive to the counterfactual demographic information given in the prompt.

### v2

@@ 3 CLIMB
We focus on the bias associated and reflected in LLMs’ diagnostic capability, where the diagnoses are identified by the International Classification of Diseases, Tenth Revision, Clinical Modification (ICD-10-CM) codes icd (2023). These well-known codes are adopted by numerous healthcare providers to catalog all medical diagnoses. It also categorizes diagnoses from high (Level 1) to low (Level 5) levels: L1 - Chapters; L2 - Blocks; L3 - Categories; L4 - Sub-categories; L5 - Full code. At L5, there are 94,739 diagnoses in total.
@@
In the medical domain, it is an important and essential issue to recognize that certain diagnoses are specific to particular demographics. This specificity arises from correct causal factors that affect different groups in varying ways. Properly acknowledging this distinction ensures accurate and fair clinical assessments, thereby improving healthcare outcomes for all demographic groups. We utilize the female-only [^3] https://www.icd10data.com/ICD10CM/Codes/Rules/Female_Diagnosis_Codes and male-only [^4] https://www.icd10data.com/ICD10CM/Codes/Rules/Male_Diagnosis_Codes codes defined in ICD-10-CM to identify sex-specific diagnoses, resulting in 3,116 and 529 diagnoses respectively. Conversely, identifying ethnicity-specific diagnoses is challenging because the absence of robust clinical evidence linking specific diagnoses to particular ethnic groups, without resorting to stereotypes or problematic assumptions, complicates the establishment of such categories Omiye et al. (2023). Therefore, we leave it as future work and only focus on the sex-specific diagnoses in this work.
@@ 3.1 Evaluation of Intrinsic Bias
Intrinsic bias studies the inherent associations between diagnosis and demographics in LLMs. In psychological research, Implicit Association Test (IAT) is a popular tool used to reveal the unconscious bias of humans by measuring the strength of subjects’ automatic associations between concepts and targets Greenwald et al. (1998). In the context of NLP, IAT has been adapted to assess the intrinsic bias of language models through representation associations between target words (e.g., names of different genders) and attribute words (e.g., terms related to careers or families) that are embedded in models’ internal representation Caliskan et al. (2016); Guo and Caliskan (2020); Bai et al. (2024). Our evaluation of intrinsic bias is also based on IAT with several important modifications.
To adapt IAT for our diagnostic task, we use diagnoses and demographic groups as attributes and targets respectively. We obtain the diagnoses from ICD-10-CM, and adopt people’s names as a proxy for different demographics following existing works Sandoval et al. (2023); An et al. (2023); Greenwald et al. (1998); Caliskan et al. (2016). These works show using names as a demographic proxy is advantageous because names inherently reflect demographic attributes like ethnicity and sex, allowing us to analyze biases without relying on explicit demographic labels, which may make the model conscious about the specific bias we are investigating (i.e., contradicted to unconsciousness of intrinsic bias).
Following Sandoval et al. (2023) and An et al. (2023), we collect the top 5 most popular names by sex (Female and Male) and ethnicity (White non-Hispanic, Black non-Hispanic, Hispanic and Asian and Pacific Islander)  [^5] We call them as “White”, ”Black”, ”Hispanic” and ”Asian” for simplicity, which is on a par with extrinsic part. from the “Popular Baby Names” dataset  [^6] https://data.cityofnewyork.us/Health/Popular-Baby-Names/25th-nujf , which provides more fine-grained ethnicity groups than traditional IAT that only provides White and Black.
@@
$\text{MAD}=\frac{1}{n}\sum_{i=1}^{n}|x_{i}-\mu|$ (1)
In embedding-based IATs such as WEAT Caliskan et al. (2016) and CEAT Guo and Caliskan (2020), association scores are calculated using embedding distances (e.g., cosine similarity) between attributes and targets that are not applicable in recent LLMs. Instead, for decoder-only LLMs that are designed for autoregressive generation, we adopt a common alternative method – probability-based bias evaluation Chen et al. (2024b); Dong et al. (2023); Nangia et al. (2020) that use next-token probability for the association calculation. Specifically, we adopt the prompt “{$d_{i}$} is related to the name:” because it clearly instructs an LLM to assign names based on diagnosis, while keeping the model unconscious about the specific bias we are investigating (i.e., ensure “unconsciousness” of intrinsic bias). We then measure the probability of each name ($n_{i}$) appearing as the next token for the association score corresponding to diagnosis $d_{i}$. For names tokenized into multiple sub-tokens, we calculate the joint probability of all sub-tokens. With the aforementioned approach, given a set of diagnoses $D=\{d_{1},d_{2},d_{3},\ldots\}$ and a set of demographic groups $G=\{g_{1},g_{2},g_{3},\ldots\}$, the association score $s\left(d_{i},g_{j}\right)$ is defined as:
$s\left(d_{i},g_{j}\right)=\frac{1}{\left|W_{g_{j}}\right|}\sum_{k=1}^{\left|W_{g_{j}}\right|}p\left(w_{k}^{g_{j}}|d_{i}\right)$ (2)
@@
$\text{\mbox{AssocMAD}}_{d_{i}}=\frac{1}{|G|}\sum_{j=1}^{|G|}\left|s\left(d_{i},g_{j}\right)-\mu_{s\left(d_{i}\right)}\right|$ (3)
@@
$\text{\mbox{AssocMAD}}_{d_{i}}=\frac{1}{|G|}\sum_{j=1}^{|G|}\left|\frac{s\left(d_{i},g_{j}\right)-\mu_{s\left(d_{i}\right)}}{\mu_{s\left(d_{i}\right)}}\right|$ (4)
@@
$\qquad\qquad\end{multlined}\text{Single-demographic AssocMAD}_{d_{i},x}\qquad\qquad\\ =\frac{1}{|G_{\text{x}}|}\sum_{j=1}^{|G_{\text{x}}|}\left|\frac{s_{\text{x}}\left(d_{i},g_{j}^{\text{x}}\right)-\mu_{s_{\text{x}}\left(d_{i}\right)}}{\mu_{s_{\text{x}}\left(d_{i}\right)}}\right|\qquad\qquad$ (5)
@@ 3.2 Evaluation of Extrinsic Bias
Extrinsic bias aims to reflect the disparity of models’ performance on end tasks towards various demographics. To set up a controlled setting and reflect the influence of the bias attribute only, we define extrinsic bias as the change in model’s performance when we switch the demographic information in a clinical note where we use the diagnosis task introduced in CliBench Ma et al. (2024) as the end task.
CliBench is the first benchmark to evaluate the diagnostic performance of recent LLMs using a large output space (i.e., ICD-10-CM). This setting makes the performance sensitive to input perturbations, allowing for a more precise examination of potential biases. It contains representative clinical tasks induced from the Medical Information Mart for Intensive Care (MIMIC)-IV database, which is comprised of de-identified Electronic Health Records (EHR) for patients admitted to the Beth Israel Deaconess Medical Center. Unlike previous medical QA datasets, the samples in CliBench are clinical admission record in real-world with detailed clinical notes, demographic information, and a large diagnosis decision space (i.e., ICD-10-CM). This allows us to reveal LLMs’ extrinsic biases in a realistic clinical setting.
From the evaluation set of CliBench, which consists of clinical cases balanced across chapters of diagnosis, service departments, and care units, we use all 199 instances with “white” ethnicity, “male” sex, and “other” insurance type 777“Medicare” and “Medicaid” are insurance programs that aim at helping minorities. “Other” covers a more privileged group of people, which is more suitable to serve as a baseline in our experiment. (i.e., most privileged group) as the default diagnosis samples for counterfactual intervention. In this evaluation set, 23.6% of the samples contain at least one sex-specific diagnosis. We consider these samples as sex-specific and others as sex-neutral. We follow the same prompt template introduced in CliBench including the demographic information and clinical notes. In Appendix A, we provide data statistics of this curated evaluation set compared to the original one in CliBench and the prompt template.
To build the counterfactual example, we replace the demographic information with a corresponding group. When the demographic information is about sex, we also replace pronouns and gendered honorifics in the instance to ensure the context remains coherent and accurate. We define extrinsic bias score as the difference in LLM’s performance on the diagnosis prediction task. In real-world clinical settings, we prioritize reducing Type 2 errors to avoid missing diagnoses. To achieve this, we use the difference in recall rate averaged on diagnostic abstraction levels while inputting default and counterfactual patient information to the models. To mitigate the potential effect of positional bias of demographic information, we computed the average score of putting demographic information before and after clinical notes. For the diagnosis prediction task, we follow CliBench to extract valid diagnosis predictions from natural language output generated by LLMs and convert the natural language to ICD code whose description has the highest cosine similarity with the text.
@@ 4.2.1 Sex-neutral Diagnoses
Table 1 reports the intrinsic bias for sex-neutral diagnoses in terms of AssocMAD, which measure the disparity across all demographic groups. A smaller value indicates a smaller disparity (i.e., less biased). We summarize several key observations from this scenario: I. Overall, all models show smaller AssocMAD at a higher level, which can be attributed to the very broad coverage of diagnoses in coarser analyses. On average, Mistral Instruct v0.2 7B and Meditron 7B (row #2 and #7) perform the best with the lowest AssocMAD. Mistral Instruct v0.2 7B achieves the best at higher levels (L1 to L3), whereas Meditron 7B achieves the best at lower levels (L4 to L5). LLaMA2 Chat 13B (#9) performs the worst across all levels, indicating the most biased. II. Interestingly, larger models such as Mistral Instruct v0.1 46.7B (#5), LLaMA2 Chat 13B (#9), and LLaMA3 Instruct 70B (#13) did not outperform their smaller 7B versions (#1, 6 and 11), which suggests that simply increasing model size does not necessarily mitigate bias and may exacerbate it due to the model’s complexity and the varied data it is exposed to during training. For the most up-to-date version of Mistral and LLaMA, Mistral v0.3 7B (#3) is more biased than LLaMA3 Instruct 8B (#11) with an averaged AssocMAD of 0.41, compared to 0.36. III. Among medically adapted models, Meditron (7B) stands out, as introduced earlier, with the lowest AssocMAD. However, with the same based model (i.e., LLaMA2 Chat 7B; #6), Asclepius 7B (#8) achieves slighter improvement. A possible reason may be that its fine-tuning data (i.e., PMC-Patients case reports) contain more relations between the diagnosis codes and patients’ demographics. IV. Models pre-trained or fine-tuned on medical corpora with incorrect association may even introduce more intrinsic bias. For example, both BioMistral DARE 7B (#4) and OpenBioLLM 8B (#12) are worse than their base models Mistral Instruct v0.1 7B (#1) and LLaMA3 Instruct 8B (#11), respectively.
We compute AssocMAD over the entire set of real-world diagnoses (i.e., ICD-10-CM), rather than relying on a sample set. Therefore, there is no need for significance testing related to diagnosis sampling. To analyze the sensitivity to name selection, we further compute AssocMAD via top-K name resampling by frequency, where $k=\{1,2,3,4\}$ and results in Table 1 are the most certain with $k=5$ (highest overall frequency). In summary, we find the above key observations remain consistent across all resamplings. We report the detailed results in the Appendix C.
@@ 4.2.2 Sex-specific Diagnoses
We report correctness of sex preference for sex-specific diagnoses in Table 3. Although some models demonstrate high correctness for either female-only or male-only diagnoses, they often fail to exhibit a balanced and reasonable sex preference for both. For instance, LLaMA2 Chat 13B shows a high correctness of 0.94 for female-only diagnoses but a low correctness of 0.14 for male-only diagnoses. This indicates it still prefers females even for male-only diagnoses. Similarly, LLaMA3 Instruct 8B shows a striking imbalance with correctness of 0.80 for female-only and 0.03 for male-only diagnoses. Although this seems to be counter-intuitive, a possible reason is their association differences are usually quite small, which suggesting model just struggles to identify these as sex-specific rather than counter-intuitively exhibiting a pronounced opposite-sex preference. This trend is also shown in the medically adapted models and is even more severe than the corresponding base version (e.g., Asclepius compared to LLaMA2 Chat 7B), which suggests that the medically adapted training may not help them learn the correct knowledge of sex preference for these sex-specific diagnoses.
@@ 4.3.1 Overall Result
We expect a model could be slightly sensitive to sex and ethnicity as they are potential causal factors to diagnosis, but non-sensitive to insurance. From Table 4, we observe: I. Decreased diagnosis performance for female instances across models. Comparing models when given factual and counterfactual sex information, we observed a consistent performance drop in almost all models when we provide counterfactual female sex information. However, when given counterfactual ethnicity or insurance, the results are mixed. II. Medical adapted LLMs are more sensitive to demographic information change. Comparing the general domain models and their corresponding medically adapted models, we observed a more significant performance change in medically adapted ones, indicating they are more sensitive to the counterfactual demographic information given in the prompt.
We notice although the results from Tables 4 do not achieve statistical significance, our pioneering benchmark provides critical evidence of extrinsic clinical bias, highlighting the fairness challenges within the clinical domain. One possible reason for the lack of significance is our deliberate minimization of changes in counterfactual interventions (i.e., altering only a few words), which helps reduce the introduction of unnecessary noise and biased assumption. However, we acknowledge that future work building on our pioneering method, could explore modifying other medical information linked with demographics (e.g., body metrics). While this could offer new insights, it also risks making biased assumptions without strict medical support, potentially distorting the detection of genuine bias and introducing contradictions in patient cases.